OpenStudy (anonymous):

Solve open parentheses 6x minus 8 close parentheses to the 1 half power equals negative 2

OpenStudy (anonymous):

(6x-8)^1/2=-2

OpenStudy (anonymous):

another way of saying to the 1 half power is sqrt

OpenStudy (anonymous):

is it 2?

OpenStudy (anonymous):

sqrt(6*x-8) = -2

OpenStudy (anonymous):

I dont think it will have a solution

OpenStudy (anonymous):

x = 2/3

OpenStudy (anonymous):

2, -2, 12, and none of the above are my options so none of the above?

OpenStudy (anonymous):

i might be wrong, i kinda suck at math but i like helping ppl

OpenStudy (anonymous):

none of the above

OpenStudy (anonymous):

none of the above

OpenStudy (anonymous):

think about it

OpenStudy (anonymous):

what number squared gives you any negative number ;)

OpenStudy (anonymous):

x*x=- number? Unle we moving into imaginary number :)

OpenStudy (anonymous):

what he said. lmao

OpenStudy (anonymous):

im still confused

OpenStudy (anonymous):

think about it

OpenStudy (anonymous):

there is nothing you can multiply by itself to get a negative number.

OpenStudy (anonymous):

so its none of the above

OpenStudy (anonymous):

sqrt(6*x-8) = -2 lets say it was sqrt(x)=-2 there is nothing you can get to make -2. Unless it is imaginary complex numbers.